Ten samples of 15 parts each were taken from an ongoing process to establish a p-chart for control. The samples and the number of defectives in each are shown in the following table:

SAMPLE n NUMBER OF DEFECTIVE ITEMS IN THE SAMPLE
1 15 3
2 15 2
3 15 2
4 15 2
5 15 0
6 15 2
7 15 1
8 15 3
9 15 2
10 15 1

a. Determine the p−p−, Sp, UCL and LCL for a p-chart of 95 percent confidence (1.96 standard deviations). (Leave no cells blank - be certain to enter "0" wherever required. Round your answers to 3 decimal places.)

p−p−
Sp
UCL
LCL

b. What comments can you make about the process?

Process is out of statistical control
Process is in statistical control
Answer #1

Solution:

Number of samples (N) = 10

Sample size (n) = 15

(a) P-bar = Number of defects / Total Number of observations

P-bar = (3 + 2 + 2 + 2 + 0 + 2 + 1 + 3 + 2 + 1) / (10 x 15)

P-bar = 18 / 150

P-bar = 0.120

Sp = SQRT [P-bar x (1 - P-bar)] / n

Sp = SQRT [0.120 x (1 - 0.120)] / 15

Sp = 0.084

UCL = P-bar + (Z x Sp)

UCL = 0.120 + (1.96 x 0.084)

UCL = 0.285

LCL = P-bar - (Z x Sp)

LCL = 0.120 - (1.96 x 0.084)

LCL = -0.045 or 0 (Number of defects cannot be negative)

LCL = 0

(b) For each of the given defects, defect rate is calculated as,

Number of defects = 1, P-bar = 1/15 = 0.067

Number of defects = 2, P-bar = 2/15 = 0.133

Number of defects = 3, P-bar = 3/15 = 0.200

Process is in statistical control, since all of the above defect rates lie within the calculated values of UCL (0.285) and LCL (0).